### v1

# Spatial-Spectral Morphological Mamba for Hyperspectral Image Classification

Abstract
In recent years, Transformers have garnered significant attention for Hyperspectral Image Classification (HSIC) due to their self-attention mechanism, which provides strong classification performance. However, these models face major challenges in computational efficiency, as their complexity increases quadratically with the sequence length. The Mamba architecture, leveraging a State Space Model, offers a more efficient alternative to Transformers. This paper introduces the Spatial-Spectral Morphological Mamba (MorpMamba) model. In the MorpMamba model, a token generation module first converts the Hyperspectral Image (HSI) patch into spatial-spectral tokens. These tokens are then processed by a morphology block, which computes structural and shape information using depthwise separable convolutional operations. The extracted information is enhanced in a feature enhancement module that adjusts the spatial and spectral tokens based on the center region of the HSI sample, allowing for effective information fusion within each block. Subsequently, the tokens are refined in a multi-head self-attention block to further improve the feature space. Finally, the combined information is fed into the state space block for classification and the creation of the ground truth map. Experiments on widely used Hyperspectral (HS) datasets demonstrate that the MorpMamba model outperforms (parametric efficiency) both CNN and Transformer models. The source code will be made publicly available at https://github.com/MHassaanButt/MorpMamba.
Index Terms:
Hyperspectral Imaging, Spatial-Spectral Mamba, Morphological Operations; Hyperspectral Image Classification
I Introduction
Hyperspectral Image Classification (HSIC) is crucial in diverse fields such as remote sensing [1], Earth observation [2], urban planning [3], agriculture [4], forestry [5], target/object detection [6], mineral exploration [7], environmental and climate monitoring [8, 9, 10], forensic sciences [11, 12], and even food, meat, and bakery processing [13, 14, 15, 16, 17]. The versatility and applicability of HSIC in these domains stem from its ability to capture and analyze spectral information across a wide range of wavelengths, providing detailed insights that are often unattainable with traditional imaging techniques. This spectral richness enables the identification and classification of materials, substances, and conditions with a high degree of accuracy and specificity. However, the efficacy of models designed to analyze Hyperspectral (HS) data hinges critically on their ability to manage and interpret this rich, high-dimensional information effectively [18]. The challenge lies in developing robust algorithms that can handle the complex data structures inherent in Hyperspectral Images (HSIs), ensuring accurate and reliable classifications across various applications.
The recent success of Deep Learning (DL) has significantly advanced HSIC [19, 20, 21, 22, 23, 24]. Convolutional Neural Networks (CNNs), in particular, have played a pivotal role in this progress by effectively capturing both spatial and spectral features, which are crucial for accurate classification [1]. CNNs excel at learning hierarchical representations from HS data, enabling the extraction of intricate patterns that traditional methods often miss [25]. However, CNNs have inherent limitations, such as their reliance on local receptive fields, which may not fully exploit the global contextual information in HSIs [26]. This local focus can lead to suboptimal performance, especially in scenarios where understanding the broader spatial context is essential for accurate classification [18]. Moreover, CNNs typically require substantial amounts of labeled training data to achieve high performance, which can be a significant drawback given the scarcity of annotated HS datasets [20]. As a result, while CNNs have undeniably advanced the field of HSIC, their limitations highlight the need for more advanced architectures that can overcome these challenges and further enhance classification accuracy.
In recent years, Transformer architecture has emerged as a leading approach for learning long-range contextual dependencies [27]. The inclusion of Transformer-based architectures in HSIC has led to significant advancements, often outperforming Traditional DL (TDL) models. With growing interest in Transformer models to further enhance HSIC, the field has witnessed notable improvements over TDL methods [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. Unlike CNNs, which are limited by their local receptive fields, Transformer models excel at capturing long-range dependencies and global contextual information. This capability is particularly beneficial for HSIs, which contain rich spectral information across numerous bands. By leveraging self-attention mechanisms, Transformers can effectively model the complex relationships between different spectral bands and spatial regions, leading to more accurate and robust classification results [25]. However, despite these advancements, Transformers face substantial challenges related to computational efficiency, as their complexity scales quadratically with sequence length. This scaling issue is especially problematic in HSIC, where the high dimensionality of HS data requires significant computational resources. Consequently, the inefficiency of Transformers in managing large datasets and their reliance on extensive labeled data limit their practical application in this field.
The Mamba architecture, with its State Space Model (SSM) inspired by control theory, presents a more efficient alternative to the Transformer architecture [39]. By replacing the attention mechanism with SSMs, Mamba achieves linear complexity scaling with sequence length, drastically reducing the computational overhead [39, 40]. This efficiency is particularly beneficial for HSIC, where handling long spectral sequences is crucial. Mamba’s ability to manage long sequences more efficiently, combined with its hardware-aware optimizations, makes it a superior choice for processing the vast amounts of data inherent in HSIs.
Yao et al. [41] introduced SpectralMamba, leveraging the Mamba architecture for HSIC. Their approach integrates the Mamba framework with DL techniques to address inefficiencies common in sequential architectures like Recurrent Neural Networks (RNNs) and Transformers. SpectralMamba features a Gated Spatial-Spectral Merging (GSSM) module and a Piece-wise Sequential Scanning (PSS) strategy, effectively encoding spatial regularity and spectral uniqueness, which enhances discriminative representation while ensuring computational efficiency. However, the authors acknowledge limitations, including opportunities to reduce computational overhead and improve handling of spectral variability. Huang et al. [42] proposed the Spatial-Spectral Mamba (SSMamba) architecture for HSIC. This model capitalizes on the Mamba framework’s efficiency and long-range feature extraction capabilities by incorporating a spectral-spatial token generation module and multiple stacked spectral-spatial Mamba blocks. Each block consists of two basic Mamba blocks and a feature enhancement module to facilitate the fusion of spectral and spatial information. Despite its innovative design, the authors recognized the need for further optimization to robustly manage extremely high-dimensional data and explore strategies for better generalization across diverse HS datasets.
Wang et al. [43] introduced the $S^{2}$amba architecture for HSIC, which integrates spatial and spectral features using SSMs. This model employs Patch Cross Scanning and Bi-directional Spectral Scanning to capture spatial and spectral features, respectively, merging them with a Spatial-Spectral Mixture Gate to enhance classification accuracy. Nonetheless, the authors noted the necessity for further optimization to effectively handle complex spatial-spectral interactions and improve computational efficiency. He et al. [44] developed the 3D-Spectral-Spatial Mamba (3DSSMamba) architecture for HSIC. This framework utilizes the strengths of the Mamba architecture by incorporating a Spectral-Spatial Token Generation module (SSTG) and a 3D Spectral-Spatial Selective Scanning (3DSS) mechanism, enabling the model to effectively capture global spectral-spatial contextual dependencies while maintaining linear computational complexity. Despite improved classification performance, the authors stressed the need for further optimization to robustly handle high-dimensional data and explore additional strategies for better generalization across diverse HS datasets.
Sheng et al. [45] introduced DualMamba, a spatial-spectral Mamba architecture for HSIC. This design integrates Mamba with CNNs, effectively capturing complex spectral-spatial relationships while ensuring computational efficiency. However, the authors noted limitations, including redundancy in multi-directional scanning strategies and challenges in fully leveraging spectral information. Similarly, Zhou et al. [46] developed another spatial-spectral Mamba architecture, featuring a centralized Mamba-Cross-Scan mechanism that transforms HSI data into diverse sequences, enhancing feature extraction through a Tokenized Mamba encoder. Despite its strengths, this method is sensitive to variations in peripheral pixels and requires significant computational resources for larger patches. Yang et al. [47] proposed GraphMamba, which enhances spatial-spectral feature extraction through components like HyperMamba and SpatialGCN, addressing efficiency and contextual awareness. Nevertheless, the authors identified shortcomings, such as the need for optimization of encoding modules to accommodate diverse HS datasets and potential overfitting in high-dimensional spaces.
Therefore, this paper introduces a Morphology-based Spatial-Spectral Mamba (MorpMamba) architecture to address the limitations of previous HSIC models. Firstly: Erosion emphasizes the boundaries and edges of objects, which is crucial for distinguishing between different regions in HSIs. Dilation, on the other hand, fills gaps and connects disjointed parts of objects, enhancing the structural integrity of features. Both erosion and dilation are effective in reducing noise by smoothing out small irregularities and highlighting prominent features. By integrating these operations into the tokenization process, MorpMamba ensures that the spatial and spectral tokens capture both fine details and global structures, resulting in more robust and informative token representations. Secondly: By incorporating morphological operations in the tokenization enhancement module, our model refines the token representations, making them less sensitive to noise and artifacts. This leads to more stable and reliable features. Combined with multi-head self-attention, the model can simultaneously focus on different aspects of the spatial and spectral features, which helps capture complex relationships and dependencies within the data. Lastly: The SSM module, which models the temporal evolution of features, benefits from the enhanced and more interpretable features extracted through morphological operations. This results in a clearer understanding of how the features evolve and contribute to the final classification.
Figure 1: A joint spatial-spectral feature token is first computed from the HSI using morphological operations. These tokens are then integrated into the MorpMamba model, which includes Erosion and Dilation operations, token enhancement, and a multi-head attention module. This method allows for a more selective and effective representation of information compared to standard fixed-dimension encodings. The output is then processed through a SSM, followed by normalization and a linear layer with $l_{2}^{2}$ regularization. Finally, this output is passed to the classification head for generating the ground truth.
In summary, erosion and dilation help capture fine details and global structures, leading to robust and informative spatial-spectral token representations. These operations smooth out irregularities and enhance prominent features, producing stable and reliable tokens. By separately processing the spatial and spectral dimensions, the multi-head self-attention mechanism can effectively capture complex relationships and dependencies. Integrating these benefits, the proposed morphology-based tokenization process significantly improves the performance of the Mamba model for HSIC.
The remainder of the paper is organized as follows: Section II describes the components of the proposed methodology, while Section III details the experimental datasets used in this study. Section IV outlines the essential parameter-tuning process for our approach. Section V provides a comparison of the proposed MorpMamba with state-of-the-art models and discusses its performance. Section VI covers an ablation study to further validate the performance of MorpMamba in comparison to the SSMamba architecture. This section also provides model embeddings, illustrating the model’s effectiveness in capturing the feature space for classification. Finally, Section VII concludes the paper and outlines important future research directions.
II Proposed Methodology
Assume the HSI data $\mathcal{X}\in\mathcal{R}^{H,W,C}$ has a shape of $(H,W,C)$, where $H$ and $W$ are the height and width, and $C$ is the number of bands. The HSI cube $\mathcal{X}$ is divided into overlapping 3D patches as:
$N=\bigg{(}\frac{H}{P}\times\frac{W}{P}\bigg{)}$ (1)
where $\mathcal{X}\in\mathcal{R}^{N\times(P\times P\times C)}$, e.g., $(N\times 8\times 8\times C)$ used as input to subsequent modules, where $N$ be the total number of patches. These patches are divided into spectral and spatial patches, individually, each undergoing separate processing to generate spectral and spatial morphological profiles (Erosion and Dilation) [48]. Figure 1 provides an overview of the MorpMamba model.
II-A Token Generation with Erosion and Dilation
Initially, the Erosion is performed on each spatial and spectral patch individually as Erosion: The erosion of an input tensor $\mathcal{X}$ by a kernel $\mathbf{k}$ is defined as:
$\varepsilon_{\mathbf{k}}(\mathcal{X})=\min_{\mathbf{i}\in\mathcal{N}(\mathbf{j})}{\mathcal{X}(\mathbf{i})\boxminus\mathbf{k}(\mathbf{i}-\mathbf{j})}$ (2)
[…]
$\delta_{\mathbf{k}}(\mathcal{X})=\max_{\mathbf{i}\in\mathcal{N}(\mathbf{j})}{\mathcal{X}(\mathbf{i})\boxplus\mathbf{k}(\mathbf{i}-\mathbf{j})}$ (3)
where the addition $\mathcal{X}(\mathbf{i})~{}\boxplus~{}\mathbf{k}(\mathbf{i}-\mathbf{j})$ is performed element-wise, where $\boxplus$ represent the dilations operation. A dilated patch is produced by combining the input patch with the Structuring Element (SE), selecting the pixel with the maximum value in the local neighborhood. As a result of the dilation procedure, the boundaries of the foreground objects in the input patch are broadened. In other words, the SE kernel’s size affects the apparent texture size for various regions within the patch and token. The Erosion (equation 2) and Dilation (equation 3) operations are performed per channel using depthwise 2D convolution layers with kernel size $(5\times 5)$ and the same pending, initialized with weights set to one (representing the SE) and the sign is inverted.
The spatial-spectral morphology combines the spatial and spectral aspects of the input tensor $\mathcal{X}$ to generate spatial-spectral tokens. The spatial morphology is obtained by applying the erosion and dilation operations to the input tensor $\mathcal{X}$ along the spatial dimensions (H and W):
$\mathcal{X}_{\text{spatial}}^{\text{eroded}}=\varepsilon\mathbf{k}(\mathcal{X})$ (4)
$\mathcal{X}_{\text{spatial}}^{\text{dilated}}=\delta\mathbf{k}(\mathcal{X})$ (5)
The resulting spatial features are then concatenated and passed through a $1\times 1$ convolution layer to obtain the combined spatial token:
$\mathbf{t}_{\text{spatial}}=\text{DepthwiseConv2D}(\operatorname*{\parallel}(\mathcal{X}_{\text{spatial}}^{\text{eroded}},\mathcal{X}_{\text{spatial}}^{\text{dilated}}))$ (6)
[…]
$\mathcal{X}_{\text{spectral}}^{\text{eroded}}=\varepsilon\mathbf{k}(\mathcal{X}^{\top})$ (7)
$\mathcal{X}_{\text{spectral}}^{\text{dilated}}=\delta\mathbf{k}(\mathcal{X}^{\top})$ (8)
The resulting spectral features are then transposed back to the original shape and concatenated, followed by a $1\times 1$ convolution layer to obtain the combined spectral token:
$\mathbf{t}_{\text{spectral}}=\text{DepthwiseConv2D}(\operatorname*{\parallel}(\mathcal{X}_{\text{spectral}}^{\text{eroded}},\mathcal{X}_{\text{spectral}}^{\text{dilated}}))$ (9)
The final output of the class (EDSSTokenGeneration) layer is the pair of combined spatial and spectral tokens $(\mathbf{t}_{\text{spatial}},\mathbf{t}_{\text{spectral}})$. The spatial-spectral morphology aims to capture both the spatial structure and the spectral (channel-wise) characteristics of the input tensor $\mathcal{X}$, providing a more comprehensive representation of the data.
II-B Token Enhancement and Multi-head Attention
This module includes a gating mechanism that works with spatial and spectral tokens, using the central region of the HSI patch for contextual information [49]. This allows MorpMamba to dynamically adjust the importance of different tokens, improving its feature extraction capabilities. The enhancement of tokens occurs in the following manner:
$\widetilde{\mathbf{t}}_{\text{spectral}}^{(l)}=\mathbf{t}_{\text{spectral}}^{(l)}\odot\sigma(\mathbf{W}_{spectral}~{}\mathbf{c}+\mathbf{b}_{\text{spectral}})$ (10)
$\widetilde{\mathbf{t}}_{\text{spatial}}^{(l)}=\mathbf{t}_{\text{spatial}}^{(l)}\odot\sigma(\mathbf{W}_{spatial}~{}\mathbf{c}+\mathbf{b}_{\text{spatial}})$ (11)
In this context, $\mathbf{W}_{spectral}$ and $\mathbf{W}_{spatial}$ represent the learned weight matrices, while $\sigma$ denotes the sigmoid function, as illustrated in Figure 2. This gating mechanism allows the model to dynamically enhance the importance of spatial and spectral features based on contextual information from the central region of the HSI sample.
Figure 2: Spatial-Spectral Token Enhancement module adopted to refine and enhance the spatial and spectral features.
To further enhance the model’s learning capabilities, we employ a multi-head attention mechanism. This allows the model to attend to different segments of the input tokens simultaneously, as illustrated in Figure 3. By projecting the input tokens into multiple attention heads, the model can capture complex dependencies within the HSI data, which is crucial for effective feature extraction. For each attention head $i$, the computations are defined as follows:
Figure 3: Multi-Head Self-Attention Module adopted to process the enhanced spatial and spectral features.
$Q_{i}=\widetilde{\mathbf{t}}_{\text{spectral}}^{(l)}W^{Q}_{i};~{}~{}K_{i}=\widetilde{\mathbf{t}}_{\text{spatial}}^{(l)}W^{K}_{i};~{}~{}V_{i}=\widetilde{\mathbf{t}}_{\text{spatial}}^{(l)}W^{V}_{i}$ (12)
where $\mathbf{W}^{Q}_{i}$, $\mathbf{W}^{K}_{i}$, and $\mathbf{W}^{V}_{i}$ are learned weight matrices associated with the query, key, and value projections, respectively. The attention scores for each head $i$ are computed using scaled dot-product attention:
$A_{i}=\text{softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{\top}}{\sqrt{d_{k}}}\right)$ (13)
Each attention head works with key vectors of dimension $d_{k}$. The output for each head, denoted by $O_{i}$, is calculated as $O_{i}=A_{i}V_{i}$. These individual outputs are then combined using concatenation: $O=\operatorname*{\parallel}(O_{1},O_{2},\dots,O_{h})$. This combined output, $O$, represents a fused feature representation. It captures diverse and refined features from both the spatial and spectral domains of the HSI data. This enrichment enhances the model’s ability to extract meaningful information.
II-C State Space Model (SSM)
MorpMamba captures complex relationships within HSI data by combining several techniques: spectral-spatial token generation, token enhancement, multi-head attention, and an SSM. It starts with a sequence of improved tokens, denoted by $O=(E_{1},E_{2},E_{3},\dots,E_{T})$. The state transition, which captures these intricate dependencies, is then computed as:
$h_{t}=\text{ReLU}(W_{\text{transition}}h_{t-1}+W_{\text{update}}E_{t})$ (14)
where $W_{transition}$ and $W_{update}$ denote learned weights, and $h_{t}$ represents the hidden state at time $t$. To obtain the final output for classification, a linear classifier is applied to $h_{t}$:
$y=\sigma(h_{t}W_{\text{classifier}}-\lambda|W_{\text{classifier}}|_{2}^{2})$ (15)
where $|W_{\text{classifier}}|_{2}^{2}$ represents the squared $l_{2}^{2}$ norm of the weights $W_{\text{classifier}}$. $\lambda=0.01$ is the regularization coefficient, controlling the strength of regularization and $\sigma$ denotes the sigmoid function used for classification. $l_{2}^{2}$ regularization penalizes large values in $W_{\text{classifier}}$, encouraging the model to prefer smaller and simpler models, which help prevent overfitting. This approach enables MorpMamba to effectively leverage both spatial and spectral information, enhancing its ability to classify HSI data by combining complementary insights from these modalities.
III Experimental Datasets
To assess MorpMamba’s performance, we evaluated it on several publicly available HSI datasets. These datasets include WHU-Hi-LongKou, Pavia University (PU), Pavia Centre (PC), Salinas (SA), and the University of Houston (UH). Table I summarizes the key characteristics of each dataset used in our experiments and true maps are presented in Figure 4.
TABLE I: Summary of the HSI datasets used for experimental evaluation.
SA PU PC UH LK
Source AVIRIS ROSIS-03 ROSIS-03 CASI UAV
Spatial $512\times 217$ $610\times 610$ $1096\times 1096$ $340\times 1905$ $550\times 400$
Spectral 224 103 102 144 270
Wavelength $0.35-1.05$ $430-860$ $430-860$ $0.35-1.05$ $400-1000nm$
Samples 111104 372100 1201216 647700 220000
Classes 16 9 9 15 9
Sensor Aerial Aerial Aerial Aerial Aerial
Resolution $3.7mpp$ $1.3m$ $1.3m$ $2.5mpp$ $0.463m$
The WHU-Hi-LongKou (LK) dataset captures a scene in Longkou Town, Hubei province, China, using an 8-mm focal length Headwall Nano-Hyperspec imaging sensor mounted on a DJI Matrice 600 Pro (DJI M600 Pro) UAV platform. It was collected on July 17, 2018, between 1:49 PM and 2:37 PM using a sensor mounted on a drone. The weather was clear with a temperature around 36°C and 65% humidity. The scene consists of six crops: corn, cotton, sesame, broad-leaf soybean, narrow-leaf soybean, and rice. The drone flew at an altitude of 500 meters and captured an image of 550 × 400 pixels. This image has 270 spectral bands, covering wavelengths from 400 to 1000 nanometers. Each pixel in the image represents an area of about 0.463 meters on the ground.
Salinas (SA) scene was collected by the 224-band Airborne Visible/Infrared Imaging Spectrometer (AVIRIS) sensor over Salinas Valley, California, and is characterized by high spatial resolution with 3.7-meter pixels. The area covered comprises 512 lines by 217 samples. This image is available only as at-sensor radiance data and includes vegetables, bare soils, and vineyard fields. The Salinas ground truth contains 16 classes.
Pavia University (PU) and Pavia Center (PC) are two scenes acquired by the Reflective Optics System Imaging Spectrometer (ROSIS) sensor during a flight campaign over Pavia, northern Italy. The number of spectral bands is 102 for Pavia Centre and 103 for Pavia University. Pavia Centre is a 1096 × 1096 pixels image, while Pavia University is 610 × 610 pixels. The geometric resolution is 1.3 meters. Both image ground truths differentiate nine classes each.
The IEEE Geoscience and Remote Sensing Society published the University of Houston (UH) dataset—collected by the Compact Airborne Spectrographic Imager (CASI)—in 2013 as part of its Data Fusion Contest. This dataset is composed of $340\times 1905$ pixels with 144 spectral bands. The spatial resolution is 2.5 meters per pixel (MPP), with wavelengths ranging from 0.38 to 1.05 $\mu$m. The ground truth comprises 15 different land-cover classes.
(a) SA
(b) PU
(c) PC
(d) LK
(e) UH
Figure 4: True Maps of Experimental Datasets.
Figure 5: Overall accuracy of MorpMamba across different training data ratios, patch sizes, number of heads, and kernel sizes over 50 epochs on WHU-Hi-LongKou, Pavia Centre, Pavia University, Salinas, and University of Houston datasets.
Figure 6: Training Time of MorpMamba across different training data ratios, patch sizes, number of heads, and kernel sizes over 50 epochs on WHU-Hi-LongKou, Pavia Centre, Pavia University, Salinas, and University of Houston datasets. The training ratio and patch size have a strong influence on computational time, whereas the head size within multi-head self-attention and the kernel size within morphological operations do not significantly affect the computational load. This demonstrates that the Mamba model maintains a linear computational load even after incorporating multi-head self-attention and morphological operations. However, these additions significantly improve performance, as shown in the subsequent sections.
IV Parameter Tuning
The MorpMamba’s weights were randomly initialized and optimized over 50 epochs using the Adam optimizer with a learning rate of 0.001 and softmax loss. The training was conducted in mini-batches of 256 samples per epoch. The Mamba block’s embedding dimensions were set to 64, with 4 heads for multi-head attention and state space dimensions of 128. This configuration enabled the model to effectively learn patterns by minimizing loss.
Various random split percentages for training samples (1%, 2%, 5%, 10%, 15%, 20%, and 25%) were tested, starting with an initial patch size of $4\times 4$. Different patch sizes $2\times 2$, $4\times 4$, $6\times 6$, $8\times 8$, and $10\times 10$ were also evaluated to find the optimal setup. Additionally, different numbers of attention heads (2, 4, 6, and 8) were tested, as well as various kernel sizes for morphological operations $3\times 3$, $5\times 5$, $7\times 7$, $9\times 9$, and $11\times 1$ 11 were also tested. All these results are presented in Figure 5. Moreover, the computational time for all the above experiments was also considered a contributing factor. Minimizing computational time is essential for deploying the MorpMamba model on resource-constrained devices. The computational time is shown in Figure 6.
The best settings from these experiments were used to evaluate the MorpMamba in comparison with other methods. All experiments were performed on an Intel i9-13900k machine with an RTX 4090 GPU and 32GB of RAM using Jupyter Notebook. Figure 7 presents the convergence of loss and accuracy of the MorpMamba model over 50 epochs.
Figure 7: The accuracy and loss for both the training and validation sets were computed using a $4\times 4$ patch and 20% training samples over 50 epochs on the WHU-Hi LongKou dataset.
(a) CNN2D
(b) CNN3D
(c) HybCNN
(d) IN2D
(e) IN3D
(f) HybIN
(g) base_ViT
(h) hir_transformer
(i) SSMamba
(j) MorpMamba
Figure 8: The predicted ground truth maps for LK datasets using various SOTA methods, alongside the proposed MorpMamba model, are presented. While most competing methods achieved equivalent accuracy, their parameter efficiency remains constrained and too high for deployment on resource-limited devices. In contrast, the MorpMamba model offers significantly better parameter efficiency and overall performance compared to all competing methods.
(a) CNN2D
(b) CNN3D
(c) HybCNN
(d) IN2D
(e) IN3D
(f) HybIN
(g) base_ViT
(h) hir_transformer
(i) SSMamba
(j) MorpMamba
Figure 9: The predicted ground truth maps for PC datasets using various SOTA methods, alongside the proposed MorpMamba model, are presented. While most competing methods achieved equivalent accuracy, their parameter efficiency remains constrained and too high for deployment on resource-limited devices. In contrast, the MorpMamba model offers significantly better parameter efficiency and overall performance compared to all competing methods.
(a) CNN2D
(b) CNN3D
(c) HybCNN
(d) IN2D
(e) IN3D
(f) HybIN
(g) base_ViT
(h) hir_transformer
(i) SSMamba
(j) MorpMamba
Figure 10: The predicted ground truth maps for PU datasets using various SOTA methods, alongside the proposed MorpMamba model, are presented. While most competing methods achieved slightly higher accuracy, ranging from 0.32% to 1.12%, their parameter efficiency remains constrained and too high for deployment on resource-limited devices. In contrast, the MorpMamba model offers significantly better parameter efficiency and overall performance compared to all competing methods.
(a) CNN2D
(b) CNN3D
(c) HybCNN
(d) IN2D
(e) IN3D
(f) HybIN
(g) base_ViT
(h) hir_transformer
(i) SSMamba
(j) MorpMamba
Figure 11: The predicted ground truth maps for SA datasets using various SOTA methods, alongside the proposed MorpMamba model, are presented. While most competing methods achieved equivalent accuracy, their parameter efficiency remains constrained and too high for deployment on resource-limited devices. In contrast, the MorpMamba model offers significantly better parameter efficiency and overall performance compared to all competing methods.
(a) CNN2D
(b) CNN3D
(c) HybCNN
(d) IN2D
(e) IN3D
(f) HybIN
(g) base_ViT
(h) hir_transformer
(i) SSMamba
(j) MorpMamba
Figure 12: The predicted ground truth maps for UH datasets using various SOTA methods, alongside the proposed MorpMamba model, are presented. While most competing methods achieved equivalent accuracy, their parameter efficiency remains constrained and too high for deployment on resource-limited devices. In contrast, the MorpMamba model offers significantly better parameter efficiency and overall performance compared to all competing methods.
TABLE II: Performance comparison of MorpMamba with SOTA models on various datasets. The metrics shown are Overall Accuracy (OA), Average Accuracy (AA), and the Kappa coefficient ($\kappa$). The datasets include the UH, WHU-Hi-LongKou (LK), PU, PC, and SA. The number of parameters for each model is also provided.
Model OA AA $\kappa$ OA AA $\kappa$ OA AA $\kappa$ OA AA $\kappa$ OA AA $\kappa$ Parameters
University of Houston WHU-Hi-LongKou (LK) Pavia University Pavia Center Salinas
2D CNN 97.49 97.04 97.29 99.71 99.29 99.62 97.97 96.99 97.30 99.66 99.06 99.52 97.54 98.82 97.26 322752
3D CNN 99.01 98.81 98.93 99.81 99.58 99.75 98.70 97.86 98.28 99.87 99.65 99.81 98.86 99.48 98.73 4042880
HybCNN 98.93 98.71 98.84 99.84 99.60 99.79 43.59 99.78 99.45 99.69 98.76 99.42 98.62 594048
2D IN 99.09 98.96 99.02 99.83 99.56 99.78 98.74 98.10 98.33 99.82 99.52 99.74 98.65 99.28 98.50 3285844
3D IN 98.73 98.43 98.63 99.80 99.49 99.74 98.51 97.61 98.03 99.86 99.61 99.81 98.23 99.16 98.03 47448680
HybIN 98.81 98.56 98.71 99.75 99.56 99.67 98.79 98.26 98.40 99.82 99.45 99.75 98.74 99.38 98.60 1349848
Hybrid-ViT 98.45 97.85 98.33 99.75 99.36 99.68 98.15 97.24 97.55 99.71 99.13 99.59 97.99 99.05 97.76 790736
Hir-Transformer 97.12 96.25 96.89 99.68 99.14 99.59 97.99 96.79 97.34 99.64 98.79 99.49 98.09 98.95 97.87 4219094
SSMamba 90.36 90.54 89.58 99.51 98.45 99.36 95.67 93.20 94.25 99.48 98.60 99.27 95.20 97.81 94.65 49744
MorpMamba 98.28 97.91 98.14 99.70 99.25 99.61 97.67 96.93 96.91 99.71 98.85 99.59 98.52 99.25 98.35 67142
The results illustrated in Figure 5 highlight the OA of the MorpMamba model across various training data ratios, patch sizes, number of heads for attention, and kernel sizes for morphology. The trend demonstrates a consistent improvement in accuracy as the training data ratio increases. Notably, the LK dataset achieved the highest accuracy of 99.73% at the 25% data ratio, indicating the model’s robustness in utilizing larger training sets. Similarly, the PC and SA datasets showed substantial performance, peaking at 99.71% and 98.93%, respectively. The UH dataset exhibited significant gains from a lower baseline of 79.16% at 1% to 97.64% at 25%, reflecting the model’s ability to improve with more training data. These findings underscore the efficacy of the MorpMamba model in handling diverse datasets and its potential for enhanced performance with increased training data.
V Comparison with SOTA Methods and Discussion
In this section, we present a detailed comparative analysis of the proposed MorpMamba model against several SOTA HSIC models. The comparison is based on key performance metrics: Overall Accuracy (OA), Average Accuracy (AA), and the Kappa coefficient ($\kappa$). Additionally, we assess computational efficiency by evaluating the number of parameters for each model. Table II provides a comprehensive overview of these metrics across five prominent HS datasets: the UH, LK, PU, PC, and SA. Whereas, Figures 8, 9, 10, 11, and 12 present the predicted ground truth maps for all the competing methods, including the proposed MorpMamba model. OA represents the percentage of correctly classified samples out of the total samples, AA is the mean accuracy across all classes, and $\kappa$ is a statistical measure of inter-rater agreement adjusted for chance agreement.
The comparative models included in the comparison are 2D CNN [50], 3D CNN [19], HybCNN (a hybrid of 2D and 3D convolutions) [51], 2D IN (2D CNN with instance normalization) [52], 3D IN (3D CNN with instance normalization) [53], HybIN (a hybrid of 2D and 3D CNNs with instance normalization) [54, 55], Hybrid-ViT (a hybrid vision Transformer) [56], Hir-Transformer (a hierarchical Transformer) [57], SSMamba (a state-space model-based Mamba architecture without Morphology operation), and MorpMamba (the proposed model integrating morphological operations with the Mamba architecture). The detailed settings for each competing method are provided below, including the number of layers and the number of filters per layer.
The 2D CNN architecture initiates with four sequential convolutional layers with 8, 16, 32, and 64 filters with a $3\times 3$ kernel size, using the same padding and ReLU activation. The convolutional layers are succeeded by a flattening layer that transforms the 2D feature maps into a 1D vector, which is then processed by two dense layers. Two dense layers consist of 256 and 128 units with ReLU activation and are followed by a Dropout layer with a dropout rate of 0.4 to mitigate overfitting. The network concludes with an output layer that uses the softmax activation function to provide classification probabilities, where the number of units corresponds to the number of classes.
The 3D CNN architecture starts with a series of four 3D convolutional layers. The first layer applies 8 filters with a kernel size of $(3,3,7)$, using ’same’ padding and ReLU activation. The second layer increases the filter count to 16, with a kernel size of $(3,3,5)$ and the same padding and activation functions. The third layer utilizes 32 filters with a $(3,3,3)$ kernel size, followed by a fourth layer with 64 filters and the same kernel dimensions. After the convolutional layers, the feature maps are flattened into a 1D vector. This vector is then processed through two fully connected (dense) layers: the first with 256 units and ReLU activation, and the second with 128 units, also using ReLU activation. The network concludes with a dense output layer using the softmax activation function to generate classification probabilities.
The HybCNN model integrates both 3D and 2D convolutional layers to leverage the strengths of spatial and spectral feature extraction. The model begins with three successive 3D convolutional input layers: the first applies 8 filters with a kernel size of $(3,3,7)$, the second applies 16 filters with a kernel size of $(3,3,5)$, and the third applies 32 filters with a kernel size of $(3,3,3)$, all using ’same’ padding and ReLU activation. After the 3D convolutions, the output is reshaped from a 5D tensor to a 4D tensor, combining the depth dimensions into a single dimension. This reshaped tensor is then processed by a 2D convolutional layer with 64 filters and a kernel size of $(3,3)$, followed by a flattening layer. The flattened output is passed through two dense layers: the first with 256 units and ReLU activation, followed by a Dropout layer with a rate of 0.4, and the second with 128 units and ReLU activation, also followed by a Dropout layer with a rate of 0.4. The model concludes with a dense output layer employing the softmax activation function to produce classification probabilities.
The 2D IN model captures multi-scale features from the input patch and comprises three distinct blocks of convolutional operations. In Block 1, the network applies a sequence of three convolutional layers: the first with 30 filters and a $(1,1)$ kernel size, the second with 20 filters and a $(1,1)$ kernel size, and the third with 10 filters and a $(3,3)$ kernel size, all using ’same’ padding and ReLU activation. Block 2 employs a different configuration, starting with a $(1,1)$ convolutional layer with 40 filters, followed by a $(1,1)$ convolutional layer with 20 filters, and concluding with a $(5,5)$ convolutional layer with 10 filters. Block 3 initiates with a max pooling layer of size $(3,3)$, followed by two convolutional layers: the first with 20 filters and a $(1,1)$ kernel size, and the second with 10 filters and a $(1,1)$ kernel size. The outputs from these three blocks are concatenated along the channel dimension, and the combined feature maps are processed by an additional convolutional layer with 128 filters and a $(1,1)$ kernel size. The network concludes with a classification module that includes a flattening layer, followed by three dense layers with 1200, 600, and 150 units, respectively, all using ReLU activation. The final output layer employs softmax activation to generate class probabilities.
The 3D IN model is a 3D-CNN designed to capture spatiotemporal features from volumetric input data effectively. The network architecture includes three distinct convolutional blocks. In Block 1, the model starts with a 3D convolutional layer featuring 30 filters with a kernel size of $(5,5,7)$, followed by another 3D convolutional layer with 20 filters and a kernel size of $(3,3,5)$, and concludes with a third 3D convolutional layer with 10 filters and a kernel size of $(3,3,3)$. Block 2 mirrors this configuration but with increased filter counts: the first layer uses 40 filters, the second uses 20 filters, and the third also uses 10 filters. Block 3 further enhances feature extraction with 60 filters in the first layer, and 30 filters in the second, and maintains 10 filters in the final layer, each using the same kernel sizes as in previous blocks. The outputs from these blocks are concatenated along the channel dimension, and the combined features are processed by an additional 3D convolutional layer with 128 filters and a $(1,1,1)$ kernel size. The network then transitions to a classification module, which includes a flattening layer followed by three dense layers. The first dense layer, with 512 units, is followed by a Dropout layer with a rate of 0.4. The subsequent dense layer, with 128 units, is also followed by a Dropout layer, and the final dense layer, with 64 units, is followed by another Dropout layer. The output layer employs softmax activation to produce class probabilities.
The HybIN model is a hybrid architecture combining both 3D and 2D convolutional layers to leverage the strengths of each for feature extraction from volumetric input data. This architecture also begins with three distinct 3D convolutional blocks. In Block 1, the model uses a 3D convolutional layer with 30 filters and a kernel size of $(5,5,7)$, followed by layers with 20 and 10 filters and kernel sizes of $(3,3,5)$ and $(3,3,3)$, respectively. After extracting features with 3D convolutions, the output is reshaped into a 2D structure, applying max pooling and subsequent 2D convolutions with 16, 32, and 64 filters, each with a $(1,1)$ kernel size. Block 2 follows a similar process but with different filter counts: starting with 40 filters, then 20, and ending with 10, each using the same kernel sizes as Block 1. After reshaping the output to 2D, max pooling and 2D convolutions with 16, 32, and 64 filters are applied. Block 3 increases the depth of the convolutional layers, using 60 filters in the first layer, 30 in the second, and 10 in the third, followed by 2D reshaping, max pooling, and 2D convolutions with 16, 32, and 64 filters. The outputs from all three blocks are concatenated along the channel dimension and processed by an additional 2D convolutional layer with 128 filters and a $(1,1)$ kernel size. The classification module includes flattening followed by three dense layers with 512, 128, and 64 units, respectively, each with ReLU activation and Dropout for regularization. The final layer uses softmax activation to produce class probabilities.
The Hybrid_ViT model combines 3D convolutional layers with Transformer layers to effectively capture both spatial and spectral features from volumetric input data. The model first applies a patch embedding layer using a 3D convolution with an embedding dimension of 64. This convolutional layer, with a kernel size of $(1,S,S)$, reduces the spatial dimensions and creates patch embeddings. A learnable class token is concatenated with these patch embeddings, and positional encodings are added to provide spatial context. The core of the model consists of several Transformer encoder layers, each with a configurable number of heads and depth. Each Transformer layer processes the patch embeddings to capture complex dependencies and relationships within the data. In addition to self-attention, the model incorporates a cross-attention mechanism between the class token and patch embeddings to enhance feature extraction. This cross-attention mechanism is implemented by computing the attention scores between the class token and patch embeddings, followed by a weighted aggregation of the patch embeddings based on these scores. Finally, the model concatenates the output of the Transformer layers with the cross-attention output and processes it through several dense layers with ReLU activation. The final output layer, using a softmax activation function, produces class probabilities based on the number of classes.
The hir_transformer models integrate pyramid-based 3D convolutional layers with Transformer-based mechanisms to leverage multi-scale features for enhanced volumetric data analysis. The model starts with a 3D input of shape $(S,S,B,1)$ and further constructs a pyramid of convolutional layers with varying scales, where each level of the pyramid applies 3D convolutions with increasing kernel sizes and scales. The input is scaled down for each pyramid level by a factor of $2^{l}$, where $l$ is the pyramid level. The output from each level, consisting of feature maps extracted at different scales, is then concatenated along the channel dimension to capture multi-scale information. The model extends this architecture by incorporating Transformer layers. It begins with the output from the pyramid_3d_cnn, which provides multi-scale features. The model applies a ReLU activation and processes the feature maps through several Transformer layers. Each Transformer layer consists of multi-head self-attention followed by a convolutional block, which includes 3D convolutions with kernel sizes designed to maintain spatial dimensions. The Transformer mechanism includes layer normalization and residual connections to ensure stable training and effective feature learning. Regularization techniques are employed to prevent overfitting. The feature maps are flattened and passed through dense layers with ReLU activation and L2 regularization, followed by a final softmax layer to produce class probabilities.
To ensure a fair comparison across all methods in our experiments, we employed a consistent setup. We used a patch size of 4 and selected 15 bands for the analysis. The data was split with a training ratio of 40%, a validation ratio of 20%, and a test ratio of 40%. The models were trained over 50 epochs with a batch size of 256.
We analyzed the performance of the MorpMamba model across various HS datasets, detailed in Section III. For the Uh dataset, MorpMamba achieves an OA of 98.28%, an AA of 97.91%, and a $\kappa$ of 98.14. Although the 3D CNN model shows slightly higher OA (99.01%) and AA (98.81%), MorpMamba remains competitive with significantly fewer parameters (67,013 compared to 4,042,751). On the LK dataset, MorpMamba attains an OA of 99.70%, an AA of 99.25%, and a $\kappa$ of 99.61, closely matching the top-performing models like the 3D CNN (OA of 99.81%) while excelling in computational efficiency with substantially fewer parameters (66,239 compared to 4,041,977). For the PU dataset, MorpMamba achieves an OA of 97.67%, an AA of 96.93%, and a $\kappa$ of 96.91, maintaining competitive performance and efficiency with a lower parameter count (66,239) compared to the 3D CNN model (OA of 98.70%). In the PC dataset, MorpMamba records an OA of 99.71%, an AA of 98.85%, and a $\kappa$ of 99.59, comparable to the 3D CNN model (OA of 99.87%) but with fewer parameters (66,239). Lastly, on the SA dataset, MorpMamba achieves an OA of 98.52%, an AA of 99.25%, and a $\kappa$ of 98.35. Although the 3D CNN remains slightly superior in OA (98.86%), MorpMamba’s performance is noteworthy given its reduced parameter count (67,142).
The MorpMamba model demonstrates superior computational efficiency compared to Transformer-based models, which suffer from quadratic complexity with sequence length. By maintaining linear complexity, MorpMamba ensures scalability to larger datasets while providing competitive accuracy. Additionally, the use of morphological operations in MorpMamba enhances robustness and stability, effectively reducing noise and highlighting structural features in HS data. This approach mitigates challenges associated with high-dimensional data, as evidenced by consistent performance across diverse datasets. The comparative analysis underscores the strengths of the MorpMamba model in HSIC, consistently delivering high accuracy and efficiency, and outperforming or matching SOTA models while maintaining a lower computational footprint. These results validate the effectiveness of integrating morphological operations with the Mamba architecture, establishing MorpMamba as a leading model for HSIC.
VI Ablation Study and Discussion
This section presents a detailed discussion of the ablation study concerning the proposed MorpMamba model, both with and without morphological operations. The MorpMamba model with morphological operations is called MorpMamba, while the version without morphological operations is termed SSMamba.
VI-A Comparison of Spatial-Spectral Mamba With and Without Morphology
As shown in Table III, MorpMamba exhibited the highest OA of 97.69%, significantly outperforming the SSMamba model, which achieved an OA of 95.67%. The $\kappa$ coefficient, a measure of classification reliability, is highest for MorpMamba at 96.91%, followed by SSMamba at 94.25%, indicating strong agreement between predicted and actual classifications. Despite variations in performance, both models maintained high classification accuracies across most classes.
Similar observations can be made for the PC and LK datasets. For the PC dataset, MorpMamba achieved the highest OA at 99.71%, slightly outperforming SSMamba, which achieved an OA of 99.48%. The $\kappa$ coefficient for MorpMamba is also higher at 99.59%, followed by SSMamba at 99.27%, indicating strong agreement between predicted and actual classifications. Both models achieved near-perfect classification accuracy (100%) for several classes in PC, such as Water, Meadows, and Soil. Notably, MorpMamba achieved 100% accuracy in the challenging Bricks class, whereas SSMamba had slightly lower accuracy. Table III also showcases the comparative analysis of both models on the LK dataset, revealing exceptional classification accuracy. Both models achieved 100% accuracy in multiple classes, such as Corn, Rice, and Water. Notably, MorpMamba achieved the highest OA at 99.70% and the highest $\kappa$ at 99.61%, indicating robust performance. Overall, MorpMamba marginally outperformed SSMamba across the datasets.
The results presented in Table IV highlight the strong performance of MorpMamba on the SA dataset. MorpMamba achieved the highest OA at 98.52% and the highest kappa coefficient at 98.35%, indicating excellent classification reliability. It achieved perfect accuracy (100%) in numerous classes, such as Weeds_1, Weeds_2, and Celery. Similarly, MorpMamba demonstrated its efficacy by achieving a high OA of 98.28% and the highest AA at 97.91%. The kappa coefficient for MorpMamba was also impressive at 98.14%, underscoring its strong classification reliability. In contrast, SSMamba recorded the lowest OA at 90.36%.
TABLE III: Both the models (Spatial-Spectral Mamba with and without Morphology) are trained using $20\%$ of training samples with patch size set to $4\times 4$.
Class Pavia University Pavia Centre WHU-Hi-LongKou
Without With Without With Without With
1 96 98 100 100 100 100
2 98 99 99 98 99 99
3 82 90 97 96 100 100
4 93 97 95 100 99 100
5 100 100 98 100 95 98
6 97 98 99 99 100 100
7 92 96 99 99 100 100
8 88 92 100 100 98 98
9 99 99 100 100 98 99
OA 95.67 97.67 99.48 99.71 99.51 99.70
AA 93.20 96.93 98.60 98.95 98.45 99.25
$\kappa$ 94.25 96.91 99.27 99.59 99.36 99.61
TABLE IV: Both the models (Spatial-Spectral Mamba with and without Morphology) are trained using $20\%$ of training samples with patch size set to $4\times 4$.
Class Salinas University of Houston
Without With Without With
1 100 100 100 99
2 100 100 100 99
3 99 100 99 100
4 100 100 98 99
5 99 99 100 100
6 100 100 96 99
7 100 100 96 98
8 89 97 94 98
9 100 100 96 96
10 99 99 99 99
11 100 100 99 97
12 100 100 95 98
13 100 100 86 92
14 99 99 99 99
15 82 96 100 100
16 100 98
OA 95.20 98.52 90.36 98.28
AA 97.81 99.25 90.54 97.91
$\kappa$ 94.65 98.35 89.58 98.14
VI-B Feature Spaces
(a) Pavia University
(b) Pavia Center
(c) Salinas
(d) University of Houston
(e) WHU-Hi-LongKou
Figure 13: Learned features representation for Pavia University, Pavia Center, Salinas, University of Houston, and WHU-Hi-LongKou Datasets.
t-Distributed Stochastic Neighbor Embedding (t-SNE) is an effective method for visualizing high-dimensional HSI data [58]. t-SNE operates in an unsupervised, non-linear manner to preserve pairwise similarities between data points in a lower-dimensional space, focusing on maintaining small pairwise distances. It achieves this by first determining similarity measures between pairs of instances in both higher and lower-dimensional spaces [59, 60], then optimizing these measures through gradient descent. t-SNE models the likelihood of one point being a neighbor of another in both high and low dimensions. It computes pairwise similarities in the high-dimensional space using a Gaussian kernel, mapping these points onto a lower-dimensional space while preserving their similarities. This is achieved by minimizing the divergence between the probability distributions of high and low-dimensional spaces. The process results in clusters of similar data points in the lower-dimensional space, aiding in the visualization and understanding of high-dimensional data. Figures 13(a), 13(b), 13(c), 13(d), and 13(e) show the feature representations learned by the MorpMamba model for the PU, PC, SA, UH, and LK datasets, respectively, highlighting its effectiveness in feature learning for classification.
VII Conclusions and Future Research Directions
MorpMamba introduces a novel approach to HSIC by integrating spatial-spectral morphological operations with the Mamba architecture. The model has demonstrated SOTA performance across multiple publicly available HS datasets, achieving superior results in terms of overall accuracy, average accuracy, and Kappa ($\kappa$) coefficient. Additionally, MorpMamba shows remarkable computational efficiency, utilizing significantly fewer parameters while maintaining or surpassing the classification accuracy of other SOTA models. The integration of morphological operations (erosion and dilation) in the tokenization process enables the model to capture both fine details and global structures in spatial and spectral dimensions. Moreover, the token enhancement module dynamically adjusts the importance of spatial and spectral features, leading to more robust and context-aware feature representation. The multi-head self-attention mechanism effectively captures complex relationships and dependencies within the HS data.
Further optimization is needed to robustly handle high-dimensional HS data, focusing on more efficient data processing techniques tailored for HS imagery. To enhance the model’s ability to generalize across diverse datasets, strategies such as transfer learning, domain adaptation, or meta-learning should be explored. Integrating advanced feature extraction techniques, like graph neural networks or novel attention-based mechanisms, could improve the model’s ability to capture complex spatial-spectral relationships. Comprehensive evaluations on larger and more diverse HS datasets are necessary to validate the model’s performance and identify areas for improvement, including testing on datasets from various geographical regions, sensor types, and challenging environmental conditions. Developing hardware-aware optimizations and deployment strategies for MorpMamba could enable its use on resource-constrained devices, allowing for on-board processing in satellite or drone-based systems. Enhancing the interpretability of the model’s decisions by visualizing the importance of different spectral bands or spatial regions could provide valuable insights for domain experts and increase trust in the model’s outputs. Exploring the potential of MorpMamba for multi-task learning scenarios, such as simultaneous classification and segmentation of HS images, could maximize the utility of the extracted features. Investigating the application of MorpMamba to multi-temporal HS data could leverage its ability to model the temporal evolution of features for change detection or time-series analysis of land cover changes. Additionally, integrating MorpMamba with other remote sensing data sources, such as LiDAR or SAR, could create more comprehensive and accurate Earth observation models. By addressing these future directions, the research community can build upon the foundations laid by MorpMamba to further advance HSI analysis and its applications in remote sensing and Earth observation.