### v2

# Spatial-Spectral Morphological Mamba for Hyperspectral Image Classification

Abstract
In recent years, the emergence of Transformers with self-attention mechanism has revolutionized the hyperspectral image (HSI) classification. However, these models face major challenges in computational efficiency, as their complexity increases quadratically with the sequence length. The Mamba architecture, leveraging a state space model (SSM), offers a more efficient alternative to Transformers. This paper introduces the Spatial-Spectral Morphological Mamba (MorpMamba) model in which, a token generation module first converts the HSI patch into spatial-spectral tokens. These tokens are then processed by morphological operations, which compute structural and shape information using depthwise separable convolutional operations. The extracted information is enhanced in a feature enhancement module that adjusts the spatial and spectral tokens based on the center region of the HSI sample, allowing for effective information fusion within each block. Subsequently, the tokens are refined through a multi-head self-attention which further improves the feature space. Finally, the combined information is fed into the state space block for classification and the creation of the ground truth map. Experiments on widely used HSI datasets demonstrate that the MorpMamba model outperforms (parametric efficiency) both CNN and Transformer models. The source code will be made publicly available at https://github.com/MHassaanButt/MorpMamba.
Index Terms:
Hyperspectral Imaging, Spatial-Spectral Mamba, Morphological Operations; Hyperspectral Image Classification
I Introduction
Hyperspectral image (HSI) classification is crucial in diverse fields such as remote sensing [1], Earth observation [2], urban planning [3], agriculture [4], forestry [5], target/object detection [6], mineral exploration [7], environmental and climate monitoring [8, 9], forensic sciences [10, 11], and even food, meat, and bakery processing [12, 13, 14, 15, 16]. The versatility and applicability of HSI classification in these domains stem from its ability to capture and analyze spectral information across a wide range of wavelengths, providing detailed insights often unattainable with traditional imaging techniques. This spectral richness enables the identification and classification of materials, substances, and conditions with high accuracy and specificity. However, the efficacy of models designed to analyze HSI data hinges critically on their ability to manage and interpret this rich, high-dimensional information effectively [17]. The challenge lies in developing robust algorithms to handle the complex data structures inherent in Hyperspectral Images (HSIs), ensuring accurate and reliable classifications across various applications.
The recent success of Deep Learning (DL) has significantly advanced HSI classification [18, 19, 20, 21, 22, 23, 24]. Convolutional Neural Networks (CNNs), in particular, have played a pivotal role in this progress by effectively capturing both spatial and spectral features, which are crucial for accurate classification [25]. CNNs excel at learning hierarchical representations from HSI data, enabling the extraction of intricate patterns that traditional methods often miss [26]. However, CNNs have inherent limitations, such as their reliance on local receptive fields, which may not fully exploit the global contextual information in HSIs [27]. This local focus can lead to suboptimal performance, especially in scenarios where understanding the broader spatial context is essential for accurate classification [17]. Moreover, CNNs typically require substantial amounts of labeled training data to achieve high performance, which can be a significant drawback given the scarcity of annotated HSI datasets [19]. As a result, while CNNs have undeniably advanced the field of HSI classification, their limitations highlight the need for more advanced architectures that can overcome these challenges and further enhance classification accuracy.
In recent years, Transformer architecture has emerged as a leading approach for learning long-range contextual dependencies [28]. The inclusion of Transformer-based architectures in HSI classification has led to significant advancements, often outperforming Traditional DL (TDL) models. With growing interest in Transformer models to further enhance HSI classification, the field has witnessed notable improvements over TDL methods [29, 30, 31, 32, 33, 34, 35, 36]. Unlike CNNs, which are limited by their local receptive fields, Transformer models excel at capturing long-range dependencies and global contextual information. This capability is particularly beneficial for HSIs, which contain rich spectral information across numerous bands. By leveraging self-attention mechanisms, Transformers can effectively model the complex relationships between different spectral bands and spatial regions, leading to more accurate and robust classification results [26]. However, despite these advancements, Transformers face substantial challenges related to computational efficiency, as their complexity scales quadratically with sequence length. This scaling issue is especially problematic in HSI classification, where the high dimensionality of HSI data requires significant computational resources. Consequently, the inefficiency of Transformers in managing large datasets and their reliance on extensive labeled data limit their practical application in this field.
The Mamba architecture, with its State Space Model (SSM) inspired by control theory, presents a more efficient alternative to the Transformer architecture [37]. By replacing the attention mechanism with SSMs, Mamba achieves linear complexity scaling with sequence length, drastically reducing the computational overhead [37, 38]. This efficiency is particularly beneficial for HSI classification, where handling long spectral sequences is crucial. Mamba’s ability to manage long sequences more efficiently, combined with its hardware-aware optimizations, makes it a superior choice for processing the vast amounts of data inherent in HSIs.
Yao et al. [39] for the first time leveraged the Mamba architecture in remote sensing for HSIC, and proposed SpectralMamba. Their approach integrates the Mamba framework with DL techniques to address inefficiencies common in sequential architectures like Recurrent Neural Networks (RNNs) and Transformers. SpectralMamba features a Gated Spatial-Spectral Merging (GSSM) module and a Piece-wise Sequential Scanning (PSS) strategy, effectively encoding spatial regularity and spectral uniqueness, which enhances discriminative representation while ensuring computational efficiency. However, the authors acknowledge limitations, including opportunities to reduce computational overhead and improve handling of spectral variability. Huang et al. [40] proposed the Spatial-Spectral Mamba (SSMamba) architecture for HSI classification. This model capitalizes on the Mamba framework’s efficiency and long-range feature extraction capabilities by incorporating a spectral-spatial token generation module and multiple stacked spectral-spatial Mamba blocks. Each block consists of two basic Mamba blocks and a feature enhancement module to facilitate the fusion of spectral and spatial information. Despite its innovative design, the authors recognized the need for further optimization to robustly manage extremely high-dimensional data and explore strategies for better generalization across diverse HSI datasets.
Wang et al. [41] introduced the $S^{2}$amba architecture for HSI classification, which integrates spatial and spectral features using SSMs. This model employs Patch Cross Scanning and Bi-directional Spectral Scanning to capture spatial and spectral features, respectively, merging them with a Spatial-Spectral Mixture Gate to enhance classification accuracy. Nonetheless, the authors noted the necessity for further optimization to effectively handle complex spatial-spectral interactions and improve computational efficiency. He et al. [42] developed the 3D-Spectral-Spatial Mamba (3DSSMamba) architecture for HSI classification. This framework utilizes the strengths of the Mamba architecture by incorporating a Spectral-Spatial Token Generation module (SSTG) and a 3D Spectral-Spatial Selective Scanning (3DSS) mechanism, enabling the model to effectively capture global spectral-spatial contextual dependencies while maintaining linear computational complexity. Despite improved classification performance, the authors stressed the need for further optimization to robustly handle high-dimensional data and explore additional strategies for better generalization across diverse HSI datasets.
Sheng et al. [43] introduced DualMamba, a spatial-spectral Mamba architecture for HSI classification. This design integrates Mamba with CNNs, effectively capturing complex spectral-spatial relationships while ensuring computational efficiency. However, the authors noted limitations, including redundancy in multi-directional scanning strategies and challenges in fully leveraging spectral information. Similarly, Zhou et al. [44] developed another spatial-spectral Mamba architecture, featuring a centralized Mamba-Cross-Scan mechanism that transforms HSI data into diverse sequences, enhancing feature extraction through a Tokenized Mamba encoder. Despite its strengths, this method is sensitive to variations in peripheral pixels and requires significant computational resources for larger patches. Yang et al. [45] proposed GraphMamba, which enhances spatial-spectral feature extraction through components like HyperMamba and SpatialGCN, addressing efficiency and contextual awareness. Nevertheless, the authors identified shortcomings, such as optimizing encoding modules to accommodate diverse HSI datasets and potential overfitting in high-dimensional spaces.
Figure 1: A joint spatial-spectral feature token is first computed from the HSI using morphological operations. These tokens are then integrated into the MorpMamba model, which includes Erosion and Dilation operations, token enhancement, and a multi-head attention module. This method allows a more selective and effective representation of information as compared to standard fixed-dimension encodings. The output is then processed through an SSM, followed by feature normalization and a linear layer with $l_{2}^{2}$ regularization. Finally, this output is passed to the classification head for generating the ground truth.
Motivated by the ability of nonlinear operations to better identify/capture the hidden patterns and internal relationships between various shapes of objects found in complex remote sensing data [46]. This paper introduces a Morphology-based Spatial-Spectral Mamba (MorpMamba) architecture to address the limitations of previous HSI classification models. Firstly: Erosion emphasizes the boundaries and edges of objects, which is crucial for distinguishing between different regions in HSIs. Dilation, however, fills gaps and connects disjointed parts of objects, enhancing the structural integrity of features. Both erosion and dilation are effective in reducing noise by smoothing out small irregularities and highlighting prominent features. By integrating these operations into the tokenization process, MorpMamba ensures that the spatial and spectral tokens capture both fine details and global structures, resulting in more robust and informative token representations. Secondly: By incorporating morphological operations in the tokenization enhancement module, our model refines the token representations, making them less sensitive to noise and artifacts. This leads to more stable and reliable features. Combined with multi-head self-attention, the model can simultaneously focus on different aspects of the spatial and spectral features, which helps capture complex relationships and dependencies within the data. Lastly: The SSM module, which models the temporal evolution of features, benefits from the enhanced and more interpretable features extracted through morphological operations. This results in a clearer understanding of how the features evolve and contribute to the final classification.
In summary, erosion and dilation help capture fine details and global structures, leading to robust and informative spatial-spectral token representations. These operations smooth out irregularities and enhance prominent features, producing stable and reliable tokens. By separately processing the spatial and spectral dimensions, the multi-head self-attention mechanism can effectively capture complex relationships and dependencies. Integrating these benefits, the proposed morphology-based tokenization process significantly improves the performance of the Mamba model for HSI classification.
The remainder of the paper is organized as follows: Section II describes the components of the proposed methodology, while Section III details the experimental datasets used in this study. Section IV outlines the essential parameter-tuning process for our approach. Section V provides a comparison of the proposed MorpMamba with state-of-the-art models and discusses its performance. Section VI covers an ablation study to further validate the performance of MorpMamba in comparison to the SSMamba architecture. This section also provides model embeddings, illustrating the model’s effectiveness in capturing the feature space for classification. Finally, Section VII concludes the paper and outlines important future research directions.
II Proposed Methodology
Assume the HSI data $\mathcal{X}\in\mathcal{R}^{H,W,C}$ has a shape of $(H,W,C)$, where $H$ and $W$ are the height and width, and $C$ is the number of bands. The HSI cube $\mathcal{X}$ is divided into overlapping 3D patches as:
$N=\bigg{(}\frac{H}{P}\times\frac{W}{P}\bigg{)}$ (1)
where $\mathcal{X}\in\mathcal{R}^{N\times(P\times P\times C)}$, e.g., $(N\times 8\times 8\times C)$ used as input to subsequent modules, where $N$ be the total number of 3D patches. These patches are divided into spectral and spatial sub-patches, individually, each undergoing separate processing to generate spectral and spatial morphological profiles (Erosion and Dilation) [47]. Figure 1 provides an overview of the MorpMamba architecture.
II-A Token Generation with Erosion and Dilation
Initially, the erosion operation is performed on each spatial and spectral patch individually and the erosion of an input tensor $\mathcal{X}$ by a kernel $\mathbf{k}$ is defined as:
$\varepsilon_{\mathbf{k}}(\mathcal{X})=\min_{\mathbf{i}\in\mathcal{N}(\mathbf{j})}{\mathcal{X}(\mathbf{i})\boxminus\mathbf{k}(\mathbf{i}-\mathbf{j})}$ (2)
[…]
$\delta_{\mathbf{k}}(\mathcal{X})=\max_{\mathbf{i}\in\mathcal{N}(\mathbf{j})}{\mathcal{X}(\mathbf{i})\boxplus\mathbf{k}(\mathbf{i}-\mathbf{j})}$ (3)
where the addition $\mathcal{X}(\mathbf{i})~{}\boxplus~{}\mathbf{k}(\mathbf{i}-\mathbf{j})$ is performed element-wise, where $\boxplus$ represent the dilations operation. A dilated patch is produced by combining the input patch with the structuring element (SE), selecting the pixel with the maximum value in the local neighborhood. As a result of the dilation operation, the boundaries of the foreground objects in the input patch are broadened. In other words, the SE kernel’s size affects the apparent texture size for various regions within the patch and token. The erosion (Equation 2) and dilation (Equation 3) operations are performed per channel using depthwise 2D convolution layers with a kernel of size $(5\times 5)$ and the same pending, initialized with weights set to one (representing the SE) and the sign is inverted.
The spatial-spectral morphology combines the spatial and spectral aspects of the input tensor $\mathcal{X}$ to generate spatial-spectral tokens. The spatial morphological feature is obtained by applying the erosion and dilation operations to the input tensor $\mathcal{X}$ along the spatial dimensions ($H$ and $W$):
$\mathcal{X}_{\text{spatial}}^{\text{eroded}}=\varepsilon\mathbf{k}(\mathcal{X})$ (4)
$\mathcal{X}_{\text{spatial}}^{\text{dilated}}=\delta\mathbf{k}(\mathcal{X})$ (5)
where the resulting spatial features are then concatenated and passed through a depthwise-convolution layer to obtain the combined spatial token:
$\mathbf{t}_{\text{spatial}}=\text{Conv2D${}_{\text{Depthwise}}$}(\operatorname*{\parallel}(\mathcal{X}_{\text{spatial}}^{\text{eroded}},\mathcal{X}_{\text{spatial}}^{\text{dilated}}),\mathcal{W})$ (6)
[…]
$\mathcal{X}_{\text{spectral}}^{\text{eroded}}=\varepsilon\mathbf{k}(\mathcal{X}^{\top})$ (7)
$\mathcal{X}_{\text{spectral}}^{\text{dilated}}=\delta\mathbf{k}(\mathcal{X}^{\top})$ (8)
where the resulting spectral features are then concatenated and passed through a depthwise-convolution layer to obtain the combined spectral token:
$\mathbf{t}_{\text{spectral}}=\text{Conv2D${}_{\text{Depthwise}}$}(\operatorname*{\parallel}(\mathcal{X}_{\text{spectral}}^{\text{eroded}},\mathcal{X}_{\text{spectral}}^{\text{dilated}}),\mathcal{W})$ (9)
where $\mathbf{t}_{\text{spectral}}\in\mathbb{R}^{N\times(P\times P\times C)}$ as the processing is similar to spatial tokens but focuses on spectral dimensions. The final output of the class (EDSSTokenGeneration) layer is the pair of combined spatial and spectral tokens $(\mathbf{t}_{\text{spatial}},\mathbf{t}_{\text{spectral}})$. The spatial-spectral morphology aims to capture both the spatial structure and the spectral (channel-wise) characteristics of the input tensor $\mathcal{X}$, providing a more comprehensive representation of the input data.
II-B Token Enhancement and Multi-head Attention
This module includes a gating mechanism that works with spatial and spectral tokens, using the central region of the HSI patch to obtain contextual information [39]. Additionally, this allows MorpMamba to dynamically adjust the importance of different tokens, improving its feature extraction capabilities. The enhancement of tokens occurs in the following manner:
$\widetilde{\mathbf{t}}_{\text{spectral}}^{(l)}=\mathbf{t}_{\text{spectral}}^{(l)}\odot\sigma(\mathbf{W}_{spectral}~{}\mathbf{c}+\mathbf{b}_{\text{spectral}})$ (10)
$\widetilde{\mathbf{t}}_{\text{spatial}}^{(l)}=\mathbf{t}_{\text{spatial}}^{(l)}\odot\sigma(\mathbf{W}_{spatial}~{}\mathbf{c}+\mathbf{b}_{\text{spatial}})$ (11)
In this context, $\mathbf{W}_{spectral}$ and $\mathbf{W}_{spatial}$ represent the learned weight matrices, while $\sigma$ denotes the sigmoid function, as illustrated in Figure 2. This gating mechanism allows the model to dynamically enhance the importance of spatial and spectral features based on contextual information extracted from the central region of the HSI sample.
Figure 2: Spatial-Spectral token enhancement module adopted to refine and enhancement of the extracted spatial and spectral features.
We employ a multi-head self-attention mechanism to further enhance the model’s learning capabilities. This allows the model to attend to different segments of the input tokens simultaneously, as illustrated in Figure 3. By projecting the input tokens into multiple attention heads, the model can capture complex dependencies within the HSI data, which is crucial for effective feature extraction during the model training. For each attention head $i$, the computations are defined as follows:
Figure 3: Multi-head self-attention module adapted to interact with the enhanced spatial and spectral features.
$Q_{i}=\widetilde{\mathbf{t}}_{\text{spectral}}^{(l)}W^{Q}_{i};~{}~{}K_{i}=\widetilde{\mathbf{t}}_{\text{spatial}}^{(l)}W^{K}_{i};~{}~{}V_{i}=\widetilde{\mathbf{t}}_{\text{spatial}}^{(l)}W^{V}_{i}$ (12)
where $\mathbf{W}^{Q}_{i}\in\mathbb{R}^{d_{\text{input}}\times d_{k}}$, $\mathbf{W}^{K}_{i}\in\mathbb{R}^{d_{\text{input}}\times d_{k}}$, and $\mathbf{W}^{V}_{i}\in\mathbb{R}^{d_{\text{input}}\times d_{v}}$ are learned weight matrices associated with the query, key, and value projections, respectively, where for each attention head, $K_{i}\in\mathbb{R}^{N\times d_{k}}$, $Q_{i}\in\mathbb{R}^{N\times d_{k}}$, $V_{i}\in\mathbb{R}^{N\times d_{v}}$ be the dimension of key, query, and value. The attention scores for each head $i$ are computed using scaled dot-product between $\mathbf{Q_{i}}$ and $\mathbf{K_{i}}$:
$A_{i}=\text{softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{\top}}{\sqrt{d_{k}}}\right)$ (13)
Each attention head works with key vectors of dimension $d_{k}$. The output for each head, denoted by $O_{i}$, is calculated as $O_{i}=A_{i}V_{i}$. These individual outputs are then combined using concatenation operations $O=\operatorname*{\parallel}(O_{1},O_{2},\dots,O_{h})$. This combined output $O$, represents a fused feature representation which captures diverse and refined features from both the spatial and spectral domains of the HSI data. This enrichment enhances the model’s ability to extract meaningful information in the context of classification.
II-C State Space Model (SSM)
MorpMamba captures complex and intrinsic relationships within HSI data by combining techniques in several stages: spectral-spatial token generation, token enhancement, multi-head self-attention, and an SSM. It starts with a sequence of enhanced tokens, denoted by $O=(E_{1},E_{2},E_{3},\dots,E_{T})$. The state transition, which captures these intricate dependencies, is then computed as:
$h_{t}=\text{ReLU}(W_{\text{trans}}h_{t-1}+W_{\text{update}}E_{t})$ (14)
where $W_{trans}$ and $W_{update}$ denote learned weights, and $h_{t}$ represents the hidden state at time $t$. To obtain the final output for classification, a linear classifier is applied to $h_{t}$:
$y=\sigma(h_{t}W_{\text{classifier}}-\lambda|W_{\text{classifier}}|_{2}^{2})$ (15)
where $|W_{\text{classifier}}|_{2}^{2}$ represents the squared $l_{2}^{2}$ norm of the weights $W_{\text{classifier}}$. $\lambda=0.01$ is the regularization coefficient, controlling the strength of regularization and $\sigma$ denotes the sigmoid function used for classification. $l_{2}^{2}$ regularization penalizes large values in $W_{\text{classifier}}$, encouraging the model to prefer smaller and simpler models, which help prevent overfitting. This approach enables MorpMamba to effectively leverage both spatial and spectral information, enhancing its ability to boost the classification performance by combining complementary insights from these modalities.
III Experimental Datasets
To assess MorpMamba’s performance, we evaluated it on several publicly available HSI datasets. These datasets include WHU-Hi-LongKou, Pavia University (PU), Pavia Centre (PC), Salinas (SA), and the University of Houston (UH). Table I summarizes the key characteristics of each dataset used in our experiments.
TABLE I: Summary of the HSI datasets used for experimental evaluation.
SA PU PC UH LK
Source AVIRIS ROSIS-03 ROSIS-03 CASI UAV
Spatial $512\times 217$ $610\times 610$ $1096\times 1096$ $340\times 1905$ $550\times 400$
Spectral 224 103 102 144 270
Wavelength $0.35-1.05$ $430-860$ $430-860$ $0.35-1.05$ $400-1000nm$
Samples 111104 372100 1201216 647700 220000
Classes 16 9 9 15 9
Sensor Aerial Aerial Aerial Aerial Aerial
Resolution $3.7mpp$ $1.3m$ $1.3m$ $2.5mpp$ $0.463m$
The WHU-Hi-LongKou (LK) dataset captures a scene in Longkou Town, Hubei province, China, using an 8-mm focal length Headwall Nano-Hyperspec imaging sensor mounted on a DJI Matrice 600 Pro (DJI M600 Pro) UAV platform. It was collected on July 17, 2018, between 1:49 PM and 2:37 PM using a sensor mounted on a drone. The weather was clear with a temperature around 36°C and 65% humidity. The scene consists of six crops: corn, cotton, sesame, broad-leaf soybean, narrow-leaf soybean, and rice. The drone flew at an altitude of 500 meters and captured an image of 550 × 400 pixels. This image has 270 spectral bands, covering wavelengths from 400 to 1000 nanometers. Each pixel in the image represents an area of about 0.463 meters on the ground.
Salinas (SA) scene was collected by the 224-band Airborne Visible/Infrared Imaging Spectrometer (AVIRIS) sensor over Salinas Valley, California, and is characterized by high spatial resolution with 3.7-meter pixels. The area covered comprises 512 lines by 217 samples. This image is available only as at-sensor radiance data and includes vegetables, bare soils, and vineyard fields. The Salinas ground truth contains 16 classes.
Pavia University (PU) and Pavia Center (PC) are two scenes acquired by the Reflective Optics System Imaging Spectrometer (ROSIS) sensor during a flight campaign over Pavia, northern Italy. The number of spectral bands is 102 for Pavia Centre and 103 for Pavia University. Pavia Centre is a 1096 × 1096 pixels image, while Pavia University is 610 × 610 pixels. The geometric resolution is 1.3 meters. Both image ground truths differentiate nine classes each.
The IEEE Geoscience and Remote Sensing Society published the University of Houston (UH) dataset—collected by the Compact Airborne Spectrographic Imager (CASI)—in 2013 as part of its Data Fusion Contest. This dataset is composed of $340\times 1905$ pixels with 144 spectral bands. The spatial resolution is 2.5 meters per pixel (MPP), with wavelengths ranging from 0.38 to 1.05 $\mu$m. The ground truth comprises 15 different land-cover classes.
Figure 4: Overall accuracy of MorpMamba across different training data ratios, patch sizes, number of heads, and kernel sizes over 50 epochs on WHU-Hi-LongKou, Pavia Centre, Pavia University, Salinas, and University of Houston datasets.
Figure 5: Training Time of MorpMamba across different training data ratios, patch sizes, number of heads, and kernel sizes over 50 epochs on WHU-Hi-LongKou, Pavia Centre, Pavia University, Salinas, and University of Houston datasets. The training ratio and patch size have a strong influence on computational time, whereas the head size within multi-head self-attention and the kernel size within morphological operations do not significantly affect the computational load.
IV Parameter Tuning
The MorpMamba’s weights were randomly initialized and optimized over 50 epochs using the Adam optimizer with a learning rate of 0.001 and softmax loss. The training was conducted in mini-batches of 256 samples per epoch. The Mamba block’s embedding dimensions were set to 64, with 4 heads for multi-head attention and state space dimensions of 128. This configuration enabled the model to effectively learn patterns by minimizing loss.
Various random split percentages for training samples (1%, 2%, 5%, 10%, 15%, 20%, and 25%) were tested, starting with an initial patch size of $4\times 4$. Different patch sizes $2\times 2$, $4\times 4$, $6\times 6$, $8\times 8$, and $10\times 10$ were also evaluated to find the optimal setup. Additionally, different numbers of attention heads (2, 4, 6, and 8) were tested, as well as various kernel sizes for morphological operations $3\times 3$, $5\times 5$, $7\times 7$, $9\times 9$, and $11\times 1$ 11 were also tested. All these results are presented in Figure 4. Moreover, the computational time for all the above experiments was also considered a contributing factor. Minimizing computational time is essential for deploying the MorpMamba model on resource-constrained devices. The computational time is shown in Figure 5.
The best settings from these experiments were used to evaluate the MorpMamba in comparison with other methods. All experiments were performed on an Intel i9-13900k machine with an RTX 4090 GPU and 32GB of RAM using Jupyter Notebook. Figure 6 presents the convergence of loss and accuracy of the MorpMamba model over 50 epochs.
Figure 6: The accuracy and loss for both the training and validation sets were computed using a $4\times 4$ patch and 20% training samples over 50 epochs on the WHU-Hi LongKou dataset.
(a) CNN2D
(b) CNN3D
(c) HybCNN
(d) IN2D
(e) IN3D
(f) HybIN
(g) Hybrid-ViT
(h) Hir-Transformer
(i) SSMamba
(j) MorpMamba
Figure 7: The predicted ground truth maps for LK datasets using various SOTA methods, alongside the proposed MorpMamba model, are presented. While most competing methods achieved equivalent accuracy, their parameter efficiency remains constrained and too high for deployment on resource-limited devices as compared to MorpMamba.
(a) CNN2D
(b) CNN3D
(c) HybCNN
(d) IN2D
(e) IN3D
(f) HybIN
(g) Hybrid-ViT
(h) Hir-Transformer
(i) SSMamba
(j) MorpMamba
Figure 8: The predicted ground truth maps for PC datasets using various SOTA methods, alongside the proposed MorpMamba model, are presented. While most competing methods achieved equivalent accuracy, their parameter efficiency remains constrained and too high for deployment on resource-limited devices as compared to MorpMamba.
(a) CNN2D
(b) CNN3D
(c) HybCNN
(d) IN2D
(e) IN3D
(f) HybIN
(g) Hybrid-ViT
(h) Hir-Transformer
(i) SSMamba
(j) MorpMamba
Figure 9: The predicted ground truth maps for PU datasets using various SOTA methods, alongside the proposed MorpMamba model, are presented. While most competing methods achieved slightly higher accuracy, ranging from 0.32% to 1.12%, their parameter efficiency remains constrained and too high for deployment on resource-limited devices as compared to MorpMamba.
(a) CNN2D
(b) CNN3D
(c) HybCNN
(d) IN2D
(e) IN3D
(f) HybIN
(g) Hybrid-ViT
(h) Hir-Transformer
(i) SSMamba
(j) MorpMamba
Figure 10: The predicted ground truth maps for SA datasets using various SOTA methods, alongside the proposed MorpMamba model, are presented. While most competing methods achieved equivalent accuracy, their parameter efficiency remains constrained and too high for deployment on resource-limited devices.
(a) CNN2D
(b) CNN3D
(c) HybCNN
(d) IN2D
(e) IN3D
(f) HybIN
(g) Hybrid-ViT
(h) Hir-Transformer
(i) SSMamba
(j) MorpMamba
Figure 11: The predicted ground truth maps for UH datasets using various SOTA methods, alongside the proposed MorpMamba model, are presented. While most competing methods achieved equivalent accuracy, their parameter efficiency remains constrained and too high for deployment on resource-limited devices.
TABLE II: Performance comparison of MorpMamba with SOTA models on various datasets. The metrics shown are Overall Accuracy (OA), Average Accuracy (AA), and the Kappa coefficient ($\kappa$). The datasets include the UH, WHU-Hi-LongKou (LK), PU, PC, and SA. The number of parameters for each model is also provided.
Model OA AA $\kappa$ OA AA $\kappa$ OA AA $\kappa$ OA AA $\kappa$ OA AA $\kappa$ Parameters
University of Houston WHU-Hi-LongKou (LK) Pavia University Pavia Center Salinas
2D CNN 97.49 97.04 97.29 99.71 99.29 99.62 97.97 96.99 97.30 99.66 99.06 99.52 97.54 98.82 97.26 322752
3D CNN 99.01 98.81 98.93 99.81 99.58 99.75 98.70 97.86 98.28 99.87 99.65 99.81 98.86 99.48 98.73 4042880
HybCNN 98.93 98.71 98.84 99.84 99.60 99.79 43.59 99.78 99.45 99.69 98.76 99.42 98.62 594048
2D IN 99.09 98.96 99.02 99.83 99.56 99.78 98.74 98.10 98.33 99.82 99.52 99.74 98.65 99.28 98.50 3285844
3D IN 98.73 98.43 98.63 99.80 99.49 99.74 98.51 97.61 98.03 99.86 99.61 99.81 98.23 99.16 98.03 47448680
HybIN 98.81 98.56 98.71 99.75 99.56 99.67 98.79 98.26 98.40 99.82 99.45 99.75 98.74 99.38 98.60 1349848
Hybrid-ViT 98.45 97.85 98.33 99.75 99.36 99.68 98.15 97.24 97.55 99.71 99.13 99.59 97.99 99.05 97.76 790736
Hir-Transformer 97.12 96.25 96.89 99.68 99.14 99.59 97.99 96.79 97.34 99.64 98.79 99.49 98.09 98.95 97.87 4219094
SSMamba 90.36 90.54 89.58 99.51 98.45 99.36 95.67 93.20 94.25 99.48 98.60 99.27 95.20 97.81 94.65 49744
MorpMamba 98.28 97.91 98.14 99.70 99.25 99.61 97.67 96.93 96.91 99.71 98.85 99.59 98.52 99.25 98.35 67142
The results illustrated in Figure 4 highlight the OA of the MorpMamba model across various training data ratios, patch sizes, number of heads for attention, and kernel sizes for morphology. The trend demonstrates a consistent improvement in accuracy as the training data ratio increases. Notably, the LK dataset achieved the highest accuracy of 99.73% at the 25% data ratio, indicating the model’s robustness in utilizing larger training sets. Similarly, the PC and SA datasets showed substantial performance, peaking at 99.71% and 98.93%, respectively. The UH dataset exhibited significant gains from a lower baseline of 79.16% at 1% to 97.64% at 25%, reflecting the model’s ability to improve with more training data. These findings underscore the efficacy of the MorpMamba model in handling diverse datasets and its potential for enhanced performance with increased training data.
V Comparison with SOTA Methods and Discussion
In this section, we present a detailed comparative analysis of the proposed MorpMamba model against several SOTA HSIC models. The comparison is based on key performance metrics: Overall Accuracy (OA), Average Accuracy (AA), and the Kappa coefficient ($\kappa$). Additionally, we assess computational efficiency by evaluating the number of parameters for each model. Table II provides a comprehensive overview of these metrics across five prominent HSI datasets: the UH, LK, PU, PC, and SA. Whereas, Figures 7, 8, 9, 10, and 11 present the predicted ground truth maps for all the competing methods, including the proposed MorpMamba model. OA represents the percentage of correctly classified samples out of the total samples, AA is the mean accuracy across all classes, and $\kappa$ is a statistical measure of inter-rater agreement adjusted for chance agreement.
The comparative models included in the comparison are 2D CNN [48], 3D CNN [18], HybCNN (a hybrid of 2D and 3D convolutions) [49], 2D IN (2D CNN with instance normalization) [50], 3D IN (3D CNN with instance normalization) [51], HybIN (a hybrid of 2D and 3D CNNs with instance normalization) [52, 53], Hybrid-ViT (a hybrid vision Transformer) [54], Hir-Transformer (a hierarchical Transformer) [55], SSMamba (a state-space model-based Mamba architecture without Morphology operation), and MorpMamba (the proposed model integrating morphological operations with the Mamba architecture). The detailed settings for each competing method are provided below, including the number of layers and the number of filters per layer.
To ensure a fair comparison across all methods in our experiments, we employed a consistent setup. We used a patch size of 4 and selected 15 bands for the analysis. The data was split with a training ratio of 40%, a validation ratio of 20%, and a test ratio of 40%. The models were trained over 50 epochs with a batch size of 256, Adam optimizer, and learning rate set to 0.001.
We analyzed the performance of the MorpMamba model across various HSI datasets, detailed in Section III. For the UH dataset, MorpMamba achieves an OA of 98.28%, an AA of 97.91%, and a $\kappa$ of 98.14. Although the 3D CNN model shows slightly higher OA (99.01%) and AA (98.81%), MorpMamba remains competitive with significantly fewer parameters (67,013 compared to 4,042,751). On the LK dataset, MorpMamba attains an OA of 99.70%, an AA of 99.25%, and a $\kappa$ of 99.61, closely matching the top-performing models like the 3D CNN (OA of 99.81%) while excelling in computational efficiency with substantially fewer parameters (66,239 compared to 4,041,977). For the PU dataset, MorpMamba achieves an OA of 97.67%, an AA of 96.93%, and a $\kappa$ of 96.91, maintaining competitive performance and efficiency with a lower parameter count (66,239) compared to the 3D CNN model (OA of 98.70%). In the PC dataset, MorpMamba records an OA of 99.71%, an AA of 98.85%, and a $\kappa$ of 99.59, comparable to the 3D CNN model (OA of 99.87%) but with fewer parameters (66,239). Lastly, on the SA dataset, MorpMamba achieves an OA of 98.52%, an AA of 99.25%, and a $\kappa$ of 98.35. Although the 3D CNN remains slightly superior in OA (98.86%), MorpMamba’s performance is noteworthy given its reduced parameter count (67,142).
The MorpMamba model demonstrates superior computational efficiency compared to Transformer-based models, which suffer from quadratic complexity with sequence length. By maintaining linear complexity, MorpMamba ensures scalability to larger datasets while providing competitive accuracy. Additionally, the use of morphological operations in MorpMamba enhances robustness and stability, effectively reducing noise and highlighting structural features in HSI data. This approach mitigates challenges associated with high-dimensional data, as evidenced by consistent performance across diverse datasets. The comparative analysis underscores the strengths of the MorpMamba model in HSIC, consistently delivering high accuracy and efficiency, and outperforming or matching SOTA models while maintaining a lower computational footprint. These results validate the effectiveness of integrating morphological operations with the Mamba architecture, establishing MorpMamba as a leading model for HSIC.
VI Ablation Study and Discussion
This section presents a detailed discussion of the ablation study concerning the proposed MorpMamba model, both with and without morphological operations. The MorpMamba model with morphological operations is called MorpMamba, while the version without morphological operations is termed SSMamba.
As shown in Table III, MorpMamba exhibited the highest OA of 97.69%, significantly outperforming the SSMamba model, which achieved an OA of 95.67%. The $\kappa$ coefficient, a measure of classification reliability, is highest for MorpMamba at 96.91%, followed by SSMamba at 94.25%, indicating strong agreement between predicted and actual classifications. Despite variations in performance, both models maintained high classification accuracies across most classes.
Similar observations can be made for the PC and LK datasets. For the PC dataset, MorpMamba achieved the highest OA at 99.71%, slightly outperforming SSMamba, which achieved an OA of 99.48%. The $\kappa$ coefficient for MorpMamba is also higher at 99.59%, followed by SSMamba at 99.27%, indicating strong agreement between predicted and actual classifications. Both models achieved near-perfect classification accuracy (100%) for several classes in PC, such as Water, Meadows, and Soil. Notably, MorpMamba achieved 100% accuracy in the challenging Bricks class, whereas SSMamba had slightly lower accuracy. Table III also showcases the comparative analysis of both models on the LK dataset, revealing exceptional classification accuracy. Both models achieved 100% accuracy in multiple classes, such as Corn, Rice, and Water. Notably, MorpMamba achieved the highest OA at 99.70% and the highest $\kappa$ at 99.61%, indicating robust performance. Overall, MorpMamba marginally outperformed SSMamba across the datasets.
The results presented in Table IV highlight the strong performance of MorpMamba on the SA dataset. MorpMamba achieved the highest OA at 98.52% and the highest kappa coefficient at 98.35%, indicating excellent classification reliability. It achieved perfect accuracy (100%) in numerous classes, such as Weeds_1, Weeds_2, and Celery. Similarly, MorpMamba demonstrated its efficacy by achieving a high OA of 98.28% and the highest AA at 97.91%. The kappa coefficient for MorpMamba was also impressive at 98.14%, underscoring its strong classification reliability. In contrast, SSMamba recorded the lowest OA at 90.36%.
TABLE III: Both the models (Spatial-Spectral Mamba with and without Morphology) are trained using $20\%$ of training samples with patch size set to $4\times 4$.
Class Pavia University Pavia Centre WHU-Hi-LongKou
Without With Without With Without With
1 96 98 100 100 100 100
2 98 99 99 98 99 99
3 82 90 97 96 100 100
4 93 97 95 100 99 100
5 100 100 98 100 95 98
6 97 98 99 99 100 100
7 92 96 99 99 100 100
8 88 92 100 100 98 98
9 99 99 100 100 98 99
OA 95.67 97.67 99.48 99.71 99.51 99.70
AA 93.20 96.93 98.60 98.95 98.45 99.25
$\kappa$ 94.25 96.91 99.27 99.59 99.36 99.61
TABLE IV: Both the models (Spatial-Spectral Mamba with and without Morphology) are trained using $20\%$ of training samples with patch size set to $4\times 4$.
Class Salinas University of Houston
Without With Without With
1 100 100 100 99
2 100 100 100 99
3 99 100 99 100
4 100 100 98 99
5 99 99 100 100
6 100 100 96 99
7 100 100 96 98
8 89 97 94 98
9 100 100 96 96
10 99 99 99 99
11 100 100 99 97
12 100 100 95 98
13 100 100 86 92
14 99 99 99 99
15 82 96 100 100
16 100 98
OA 95.20 98.52 90.36 98.28
AA 97.81 99.25 90.54 97.91
$\kappa$ 94.65 98.35 89.58 98.14
(a) Pavia University
(b) Pavia Center
(c) Salinas
(d) University of Houston
(e) WHU-Hi-LongKou
Figure 12: Learned features representation for Pavia University, Pavia Center, Salinas, University of Houston, and WHU-Hi-LongKou Datasets.
t-Distributed Stochastic Neighbor Embedding (t-SNE) is an effective method for visualizing high-dimensional HSI data. t-SNE operates in an unsupervised, non-linear manner to preserve pairwise similarities between data points in a lower-dimensional space, focusing on maintaining small pairwise distances. It achieves this by first determining similarity measures between pairs of instances in both higher and lower-dimensional spaces [56], then optimizing these measures through gradient descent. t-SNE models the likelihood of one point being a neighbor of another in both high and low dimensions. It computes pairwise similarities in the high-dimensional space using a Gaussian kernel, mapping these points onto a lower-dimensional space while preserving their similarities. This is achieved by minimizing the divergence between the probability distributions of high and low-dimensional spaces. The process results in clusters of similar data points in the lower-dimensional space, aiding in the visualization and understanding of high-dimensional data. Figures 12(a), 12(b), 12(c), 12(d), and 12(e) show the feature representations learned by the MorpMamba model for the PU, PC, SA, UH, and LK datasets, respectively, highlighting its effectiveness in feature learning for classification.
VII Conclusions and Future Research Directions
MorpMamba introduces a novel approach to HSI classification by integrating spatial-spectral morphological operations with the Mamba architecture. The model has demonstrated SOTA performance across multiple publicly available HSI datasets, achieving superior results in terms of overall accuracy, average accuracy, and Kappa ($\kappa$) coefficient. Additionally, MorpMamba shows remarkable computational efficiency, utilizing significantly fewer parameters while maintaining or surpassing the classification accuracy of other SOTA models. The integration of morphological operations (erosion and dilation) in the tokenization process enables the model to capture both fine details and global structures in spatial and spectral dimensions. Moreover, the token enhancement module dynamically adjusts the importance of spatial and spectral features, leading to more robust and context-aware feature representation. The multi-head self-attention mechanism effectively captures complex relationships and dependencies within the HSI data.
Further optimization is needed to robustly handle high-dimensional HSI data, focusing on more efficient data processing techniques tailored for HSI imagery. To enhance the model’s ability to generalize across diverse datasets, strategies such as transfer learning, domain adaptation, or meta-learning should be explored. Integrating advanced feature extraction techniques, like graph neural networks or novel attention-based mechanisms, could improve the model’s ability to capture complex spatial-spectral relationships. Comprehensive evaluations on larger and more diverse HSI datasets are necessary to validate the model’s performance and identify areas for improvement, including testing on datasets from various geographical regions, sensor types, and challenging environmental conditions. Developing hardware-aware optimizations and deployment strategies for MorpMamba could enable its use on resource-constrained devices, allowing for on-board processing in satellite or drone-based systems. Enhancing the interpretability of the model’s decisions by visualizing the importance of different spectral bands or spatial regions could provide valuable insights for domain experts and increase trust in the model’s outputs. Exploring the potential of MorpMamba for multi-task learning scenarios, such as simultaneous classification and segmentation of HSI images, could maximize the utility of the extracted features. Investigating the application of MorpMamba to multi-temporal HSI data could leverage its ability to model the temporal evolution of features for change detection or time-series analysis of land cover changes. Additionally, integrating MorpMamba with other remote sensing data sources, such as LiDAR or SAR, could create more comprehensive and accurate Earth observation models. By addressing these future directions, the research community can build upon the foundations laid by MorpMamba to further advance HSI analysis and its applications in remote sensing and Earth observation.
Acknowledgment
The authors extend their appreciation to the Researchers Supporting Project number (RSPD2024R848), King Saud University, Riyadh, Saudi Arabia.